# The Nesterov–Spokoiny Acceleration Achieves Strict o(1/k2) Convergence

### v4

@@ 6 Experiments
We compare the NSA algorithm (Algorithm 3) with classic gradient descent and benchmark acceleration methods, including the AFBM method Nesterov, (1983, 2013); Attouch and Peypouquet, (2016), the FISTA algorithm Beck and Teboulle, (2009), and the NSA algorithm in its original form Nesterov, (2017). It has been empirically observed that NAG with over-damping may converge faster (Su et al., , 2016). For this reason, we compare NSA with possibly over-damped acceleration methods. Typically, the damping factor $\geq 3$ yields better empirical performance. In all experiments, we pick the same values of learning rate $\eta$ for all methods.
The start point $\mathbf{x}_{0}$ is also same for all algorithms. In the first series of experiments (Figure 2), we assess the methods across various standard machine learning and optimization tasks, such as regression, logistic regression, and matrix completion. In the second series of experiments (Figure 3), we examine the methods in the context of standard fully-connected neural network training using the iris dataset Fisher, (1936).
@@ 1. Figure 2a:
    where $A$ is a full-rank $400\times 200$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\>\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $400\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 2. Figure 2b:
    where $A=\left(A_{1},...,A_{200}\right)^{\top}$ is a $200\times 5$ random matrix sampled from $i.i.d.\>\mathcal{N}(0,1)$ and $\mathbf{y}=(y_{1},\cdots,y_{200})^{\top}$ is a $200\times 1$ random vector of labels sampled from $i.i.d.\>Ber(1,0.5)$. We set $\eta=0.005$ for all algorithms.
@@ 3. Figure 2c:
    where $A$ is a full-rank $400\times 200$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\>\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $400\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 5. Figure 2e:
    where $A$ is a full-rank $400\times 200$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\>\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $400\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 6. Figure 2f:
    where $A_{ob}$ includes the available elements in matrix $A$ and only 20% of the elements in $A$ can be observed. The position of these elements is randomly chosen and the original matrix $A$ is a rank-3 $50\times 40$ matrix with the eigenvalues —$\left[1,2,3\right]$. To generate $A$, we generate $U,S$ and $V$ where $S=diag\{1,2,3\}$ and $U,V^{\top}$ are rank-3 $50\times 7$ matrices generated from $i.i.d\>Uniform(0,1)$. Then we calculate $A$ by $A=USV$, and set $\lambda=0.05$. We set $\eta=1$ for all algorithms.
Figure 2: Comparison of Algorithm 3 (referred to as ‘New NSA’) with benchmark methods. In the graph legend, GD represents Gradient Descent; NSA denotes the original acceleration method by Nesterov and Spokoiny Nesterov, (2017); and FISTA and AFBM methods were introduced in Beck and Teboulle, (2009) and Nesterov, (1983, 2013), respectively. In this context, $p$ is used to indicate the damping factor. As the graphs for NSA and AFBM are closely aligned when using the same damping factor, we present only the NSA method with a damping factor of $p=3$ and the AFBM method with a damping factor of $p=4$.
@@ 2. Figure 3b:
Figure 3: Comparison of Algorithm 3 (referred to as ‘New NSA’) with benchmark methods. In the graph legend, GD represents Gradient Descent; NSA denotes the original acceleration method by Nesterov and Spokoiny Nesterov, (2017); and FISTA and AFBM methods were introduced in Beck and Teboulle, (2009) and Nesterov, (1983, 2013), respectively. In this context, $p$ is used to indicate the damping factor. We present only the NSA method with a damping factor of $p=5$ and the AFBM method with a damping factor of $p=6$.
7 Proofs of Theorem 3 and Corollary 1
@@
where $[\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{k}]=\mathbf{V}\in$$\{\mathbf{X}\in\mathbb{R}^{n\times k}:\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{k}\}$ is uniformly sampled.
Select $k=n$, then $[\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{n}]$ is an orthonormal bases for $\mathbb{R}^{n}$, and we have
@@
Let $v_{i}^{k}$ represents the $k$-th element of $\mathbf{v_{i}}$ Using the Taylor formula,we have
@@
$[\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{k}]=\mathbf{V}\in$$\{\mathbf{X}\in\mathbb{R}^{n\times n}:\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{n}\}$ So it is obvious that $\mathbf{V}$ is an orthogonal matrix and $\sum_{i=1}^{n}v_{i}^{k}v_{i}^{p}$ is the inner product between the $k$-th column and the $p$-th column. According to the proposition of the orthogonal matrix, we know $\sum_{i=1}^{n}v_{i}^{k}v_{i}^{p}=\delta_{kp}$ , where $\delta_{kp}$ is the Kronecker delta. So we can simplify the formula above as
@@
8 Conclusion
This paper studies the NSA algorithm – a variant of an acceleration method of Nesterov and Spokoiny. NSA converges at rate $o(k^{-2})$ for general smooth convex programs. Also, we show that in a continuous-time perspective, the NSA algorithm is governed by a system of first-order ODEs.
Appendix A Standard Conventions for Smooth Convex Optimization
To be self-contained, we briefly review some common conventions for smooth convex optimization. In his book Nesterov, (2018), Nesterov’s discussion on smooth convex optimization focuses on a class of functions $\mathscr{F}$ such that, for any $f\in\mathscr{F}$, $\nabla f(\mathbf{x}^{*})=0$ implies that $\mathbf{x}^{*}$ is a global minimizer of $f$. In addition, $\mathscr{F}$ contains linear functions, and is closed under addition and nonnegative scaling.
Following Nesterov’s convention Nesterov, (2018), we write $f\in\mathscr{F}_{L}^{h,l}(\mathbb{R}^{n})$ ($l\leq h$) if $f\in\mathscr{F}$, and in addition,
  • $f$ is convex;
  • $f$ is $h$-times continuously differentiable;
Appendix B Additional Experiments
B.1 Zeroth-order version of NSA
We compare the NSA-zero algorithm (2) with classic gradient descent and the Nesterov Accelerated Gradient (NAG) Nesterov, (1983). In all experiments, we pick the same values of $p$ and $\eta$ for both NSA-zero and NAG. The start point $\mathbf{x}_{0}$ is also same for both NSA-zero and NAG. We choose the gradient estimation $\mathcal{G}$ as
@@ 1. Figure 4a:
    where $A$ is a $200\times 100$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\>\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $200\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 2. Figure 4b:
    where $A=\left(A_{1},...,A_{200}\right)^{\top}$ is a $200\times 5$ random matrix sampled from $i.i.d.\>\mathcal{N}(0,1)$ and $\mathbf{y}=(y_{1},\cdots,y_{200})^{\top}$ is a $200\times 1$ random vector of labels sampled from $i.i.d.\>Ber(1,0.5)$. We set $\eta=0.0005$ for all algorithms.
@@ 3. Figure 4c:
    where $A$ is a $200\times 100$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\>\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $200\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 5. Figure 4e:
    where $A$ is a $200\times 100$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\>\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $200\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 6. Figure 4f:
    where $A_{ob}$ includes the available elements in matrix $A$ and only 20% of the elements in $A$ can be observed. The position of these elements is randomly chosen and the original matrix $A$ is a rank-5 $10\times 8$ matrix with the eigenvalues —$\left[1,2,3,4,5\right]$. To generate $A$, we generate $U,S$ and $V$ where $S=diag\{1,2,3,4,5\}$ and $U,V^{\top}$ are rank-3 $50\times 7$ matrices generated from $i.i.d\>Uniform(0,1)$. Then we calculate $A$ by $A=USV$, and set $\lambda=0.005$. We set $\eta=0.05$ for all algorithms.
Figure 4: Comparison of Algorithm 2 (referred to as ‘New NSA’) with benchmark methods. In the graph legend, GD represents Gradient Descent; NSA denotes the original acceleration method by Nesterov and Spokoiny Nesterov, (2017); and FISTA and AFBM methods were introduced in Beck and Teboulle, (2009) and Nesterov, (1983, 2013), respectively. In this context, $p$ is used to indicate the damping factor. As the graphs for NSA and AFBM are closely aligned when using the same damping factor, we present only the NSA method with a damping factor of $p=3$ and the AFBM method with a damping factor of $p=4$. In this experiments, all methods use inexact gradient obtained from zeroth-order information.

### v5

@@ 6 Experiments
We compare the NSA algorithm (Algorithm 3) with classic gradient descent and benchmark acceleration methods, including the AFBM method [nesterov1983method, Nesterov2013, doi:10.1137/15M1046095], the FISTA algorithm [2009A], and the NSA algorithm in its original form [2017Random]. It has been empirically observed that NAG with over-damping may converge faster [JMLR:v17:15-084]. For this reason, we compare NSA with possibly over-damped acceleration methods. Typically, the damping factor $\geq 3$ yields better empirical performance. In all experiments, we pick the same values of learning rate $\eta$ for all methods.
The start point $\mathbf{x}_{0}$ is also same for all algorithms. In the first series of experiments (Figure 2), we assess the methods across various standard machine learning and optimization tasks, such as regression, logistic regression, and matrix completion. In the second series of experiments (Figure 3), we demonstrate the benefits of these methods by applying them to a standard fully-connected neural network trained on the Iris dataset [iris_53], highlighting their utility in key real-world applications.
@@ 1. Figure 2a:
    where $A$ is a full-rank $400\times 200$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $400\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 2. Figure 2b:
    where $A=\left(A_{1},...,A_{200}\right)^{\top}$ is a $200\times 5$ random matrix sampled from $i.i.d.\>\mathcal{N}(0,1)$ and $\mathbf{y}=(y_{1},\cdots,y_{200})^{\top}$ is a $200\times 1$ random vector of labels sampled from i.i.d.Ber(1,0.5)i.i.d.\>Ber(1,0.5). We set $\eta=0.005$ for all algorithms.
@@ 3. Figure 2c:
    where $A$ is a full-rank $400\times 200$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $400\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 5. Figure 2e:
    where $A$ is a full-rank $400\times 200$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $400\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 6. Figure 2f:
    where $A_{ob}$ includes the available elements in matrix $A$ and only 20% of the elements in $A$ can be observed. The position of these elements is randomly chosen and the original matrix $A$ is a rank-3 $50\times 40$ matrix with the eigenvalues $\left[1,2,3\right]$. To generate $A$, we generate $U,S$ and $V$ where $S=diag\{1,2,3\}$ and $U,V^{\top}$ are rank-3 $50\times 7$ matrices generated from i.i.dUniform(0,1)i.i.d\>Uniform(0,1). Then we calculate $A$ by $A=USV$, and set $\lambda=0.05$. We set $\eta=1$ for all algorithms.
Figure 2: Comparison of Algorithm 3 (referred to as ‘New NSA’) with benchmark methods. In the graph legend, GD represents Gradient Descent; NSA denotes the original acceleration method by Nesterov and Spokoiny [2017Random]; and FISTA and AFBM methods were introduced in [2009A] and [nesterov1983method, Nesterov2013], respectively. In this context, $p$ is used to indicate the damping factor. As the graphs for NSA and AFBM are closely aligned when using the same damping factor, we present only the NSA method with a damping factor of $p=3$ and the AFBM method with a damping factor of $p=4$.
@@ 2. Figure 3b:
Figure 3: Comparison of Algorithm 3 (referred to as ‘New NSA’) with benchmark methods ($y$-axis in logarithmic scale). In the graph legend, GD represents Gradient Descent; NSA denotes the original acceleration method by Nesterov and Spokoiny [2017Random]; and FISTA and AFBM methods were introduced in [2009A] and [nesterov1983method, Nesterov2013], respectively. In this context, $p$ is used to indicate the damping factor. We present only the NSA method with a damping factor of $p=5$ and the AFBM method with a damping factor of $p=6$.
7 Conclusion
This paper studies the NSA algorithm – a variant of an acceleration method of Nesterov and Spokoiny. NSA converges at rate $o(k^{-2})$ for general smooth convex programs. To our knowledge, this work is the first to establish an $o(k^{-2})$ convergence rate for the function value and an $o(\frac{1}{k^{3}\log k})$ rate for the squared gradient norm, all while maintaining strictly monotonic descent of function value.
This work also presents a comprehensive study of the NSA algorithm, providing a zeroth-order variant capable of handling inexact gradients, a extension that handles convex composite problems, and a continuous-time analysis of its dynamics.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT (Chat-4o-mini) in order to improve readability and polish writing. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendix A Standard Conventions for Smooth Convex Optimization
To be self-contained, we briefly review some common conventions for smooth convex optimization. In his book [nest2018], Nesterov’s discussion on smooth convex optimization focuses on a class of functions $\mathscr{F}$ such that, for any $f\in\mathscr{F}$, $\nabla f(\mathbf{x}^{*})=0$ implies that $\mathbf{x}^{*}$ is a global minimizer of $f$. In addition, $\mathscr{F}$ contains linear functions, and is closed under addition and nonnegative scaling.
Following Nesterov’s convention [nest2018], we write $f\in\mathscr{F}_{L}^{h,l}(\mathbb{R}^{n})$ ($l\leq h$) if $f\in\mathscr{F}$, and in addition,
  • $f$ is convex;
  • $f$ is $h$-times continuously differentiable;
Appendix B Proofs of Theorem 3 and Corollary 1
@@
where $\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{k}]=\mathbf{V}\in$ $\{\mathbf{X}\in\mathbb{R}^{n\times k}:\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{k}\}$ is uniformly sampled.
Select $k=n$, then $\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{n}]$ is an orthonormal bases for $\mathbb{R}^{n}$, and we have
@@
Let $v_{i}^{k}$ represents the $k$-th element of $\mathbf{v_{i}}$ . Using the Taylor formula,we have
@@
$\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{k}]=\mathbf{V}\in$ $\{\mathbf{X}\in\mathbb{R}^{n\times n}:\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{n}\}$ So it is obvious that $\mathbf{V}$ is an orthogonal matrix and $\sum_{i=1}^{n}v_{i}^{k}v_{i}^{p}$ is the inner product between the $k$-th column and the $p$-th column. According to the proposition of the orthogonal matrix, we know $\sum_{i=1}^{n}v_{i}^{k}v_{i}^{p}=\delta_{kp}$ , where $\delta_{kp}$ is the Kronecker delta. So we can simplify the formula above as
@@
Appendix C Additional Experiments
C.1 Zeroth-order version of NSA
We compare the NSA-zero algorithm (2) with classic gradient descent and the Nesterov Accelerated Gradient (NAG) [nesterov1983method]. In all experiments, we pick the same values of $p$ and $\eta$ for both NSA-zero and NAG. The start point $\mathbf{x}_{0}$ is also same for both NSA-zero and NAG. We choose the gradient estimation $\mathcal{G}$ as
@@ 1. Figure 4a:
    where $A$ is a $200\times 100$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $200\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 2. Figure 4b:
    where $A=\left(A_{1},...,A_{200}\right)^{\top}$ is a $200\times 5$ random matrix sampled from $i.i.d.\>\mathcal{N}(0,1)$ and $\mathbf{y}=(y_{1},\cdots,y_{200})^{\top}$ is a $200\times 1$ random vector of labels sampled from i.i.d.Ber(1,0.5)i.i.d.\>Ber(1,0.5). We set $\eta=0.0005$ for all algorithms.
@@ 3. Figure 4c:
    where $A$ is a $200\times 100$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $200\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 5. Figure 4e:
    where $A$ is a $200\times 100$ random matrix with entries sampled as $i.i.d.$ standard Gaussian, $\mathcal{N}(0,1)$ and $\mathbf{b}$ is a $200\times 1$ standard Gaussian vector. We set $\eta=0.0005$ for all algorithms.
@@ 6. Figure 4f:
    where $A_{ob}$ includes the available elements in matrix $A$ and only 20% of the elements in $A$ can be observed. The position of these elements is randomly chosen and the original matrix $A$ is a rank-5 $10\times 8$ matrix with the eigenvalues $\left[1,2,3,4,5\right]$. To generate $A$, we generate $U,S$ and $V$ where $S=diag\{1,2,3,4,5\}$ and $U,V^{\top}$ are rank-3 $50\times 7$ matrices generated from $i.i.d$ $\mathrm{Uniform}(0,1)$. Then we calculate $A$ by $A=USV$, and set $\lambda=0.005$. We set $\eta=0.05$ for all algorithms.
Figure 4: Comparison of Algorithm 2 (referred to as ‘New NSA’) with benchmark methods. In the graph legend, GD represents Gradient Descent; NSA denotes the original acceleration method by Nesterov and Spokoiny [2017Random]; and FISTA and AFBM methods were introduced in [2009A] and [nesterov1983method, Nesterov2013], respectively. In this context, $p$ is used to indicate the damping factor. As the graphs for NSA and AFBM are closely aligned when using the same damping factor, we present only the NSA method with a damping factor of $p=3$ and the AFBM method with a damping factor of $p=4$. In this experiments, all methods use inexact gradient obtained from zeroth-order information.